# Deep Metric Imitation Learning for Stable Motion Primitives

### v2

@@ 4.2.3 Loss Function
Notably, we can enforce the surrogate stability conditions from Theo. 4 within a DNN by minimizing the following expression $\forall y_{0}\in\mathcal{L}$ and $\forall t\in\mathbb{R}_{\geq 0}$:
@@
where $m\in\mathbb{R}_{>0}$. This function resembles the form of the triplet loss introduced in Sec. 3.4. In this context, $y_{\mathrm{g}}$ serves as the anchor sample, $y_{t+\Delta t}$ as the positive sample, and $y_{t}$ as the negative sample. Let us explore how this expression induces the fulfillment of the surrogate stability conditions within a DNN.
@@
Figure 5: Left: Illustration of the effect of optimizing $\ell_{\text{stable}}$ in one triplet. Right: Example of trajectories generated with $f^{\mathcal{T}}_{\theta}$ and $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$ post-training. Each $y_{t+\Delta t}$ is closer to $y_{\mathrm{g}}$ than its predecessor $y_{t}$.
For any $y_{0}$ where $x_{0}\neq x_{\mathrm{g}}$, minimizing (16) enforces transitions to progressively approach $y_{\mathrm{g}}$. This is achieved as it implies that
$d(y_{\mathrm{g}},y_{t+\Delta t})+m\leq d(y_{\mathrm{g}},y_{t}),$ (17)
and, therefore, $d(y_{\mathrm{g}},y_{t+\Delta t})<d(y_{\mathrm{g}},y_{t})$, i.e., the second surrogate condition (see Fig. 5).
@@
In contrast, the state $x_{\mathrm{g}}$ maps to $y_{\mathrm{g}}$ by construction, which modifies the behavior of (16) when $y_{0}=y_{\mathrm{g}}=\psi_{\theta}(x_{\mathrm{g}})$. In this case, for $t=0$, the distance $d(y_{\mathrm{g}},y_{0})$ cannot be modified through the parameters of the DNN, since it is zero by construction. Hence, the minimum of (16) is achieved at $m$, provided that $y_{0+\Delta t}$ also equals $y_{\mathrm{g}}$. Moreover, under the same logic, for this initial condition, any $y_{t+\Delta t}$ gets modified towards the value of $y_{\mathrm{g}}$. Consequently, the first surrogate condition is enforced.
@@
In this equation, $\mathcal{B}^{s}$ denotes a batch of initial latent states $y_{0}$. These states are derived by mapping initial states $x_{0}$ (sampled randomly from $\mathcal{T}$) via the function $\psi_{\theta}$. Meanwhile, $\mathcal{H}^{s}$ represents a set of time points $t$ at which the loss is minimized.
@@ 4.3 On the Stability Loss Metric
To elaborate on this point, recall that the learned dynamical system can be expressed as $\dot{x}_{t}=\phi_{\theta}(y_{t})$. It then follows that the output of our model is entirely determined by the latent state $y_{t}$. This implies that if two different states $x^{\text{a}}_{t}$ and $x^{\text{b}}_{t}$ map to the same latent state, that is, $y_{t}=\psi_{\theta}(x^{\text{a}}_{t})=\psi_{\theta}(x^{\text{b}}_{t})$, the time derivative computed by the learned dynamical systems for both states will be identical. Such behavior would manifest in certain states if $\psi_{\theta}$ were a non-bijective [^4] More rigorously, the property being described is that of non-injective functions. However, for practical purposes, we can define the codomain of $\psi_{\theta}$ to be equal to its image, which is the property that an injective function must have to be bijective. Hence, in this specific context, we use these terms interchangeably. function. It would, therefore, be problematic if the learning process were to force $\psi_{\theta}$ to be non-bijective, as this would constrain the family of solutions to which the system can converge, harming the DNN’s optimization process. Ideally, we would like $\psi_{\theta}$ to have the capacity to converge to a bijective function, such that every state $x_{t}$ maps to a unique latent state $y_{t}$. Consequently, for any state $x_{t}$, it would be possible to compute a value of $\dot{x}_{t}$ that is independent of those calculated for any other state. Fig. 6 presents an example of bijective and non-bijective mappings between dynamical systems.
@@ 4.3.1 Homeomorphisms and State Space Geometry
In our context, this implies that the distance function employed in the stability loss (18) must induce a topology in $\mathcal{L}$ that is equivalent to that of $\mathcal{T}$. For example, if orientations are described using unit quaternions, the topology of $\mathcal{T}$ would be spherical. Then, the stability loss metric should generate a topology that is homeomorphic to the sphere. Otherwise, it would be infeasible for the DNN to establish a homeomorphism between $\mathcal{T}$ and $\mathcal{L}$. In Sec. 5, we utilize unit quaternions to represent orientations; for a detailed discussion on metrics and pose control in this context, the reader is referred to Appendix B.
@@ 4.4 Boundary Conditions
Lastly, it is crucial to ensure a dynamical system evolving in $\mathcal{T}$ always remains within this manifold. Two scenarios are relevant for this work: 1) $\mathcal{T}$ is not positively invariant with respect to $f^{\mathcal{T}}_{\theta}$, and 2) Euclidean operations are employed in non-Euclidean state spaces.
@@ 4.4.1 Positively Invariant Sets
To address this scenario, we design the dynamical system so that it cannot leave $\mathcal{T}$ by construction. This can be achieved by projecting any transitions that would leave $\mathcal{T}$ back onto its boundary. For example, in Euclidean state spaces, $\mathcal{T}$ can be represented as a hypercube; therefore, in this case, this projection is achieved by saturating/clipping the points that would leave $\mathcal{T}$. Furthermore, we have found that introducing an additional loss, denoted as $\ell_{\partial}$, for optimizing the DNN can aid the optimization process. As proposed in [6], this can be accomplished by calculating the scalar product between the dynamical system’s velocity $v(x_{t})$ (which equates to $f^{\mathcal{T}}_{\theta}$ for first-order systems) and the outward-pointing normal vector $n(x_{t})$ at states within the boundary of $\mathcal{T}$. This product should be ensured to be equal to or less than zero. For DNNs, this can be achieved by minimizing
@@
4.4.2 Non-Euclidean State Spaces
@@
Non-Euclidean state spaces are commonly defined within a higher-dimensional Euclidean space and can sometimes be constructed by incorporating constraints into this space. Consider the unit quaternion as an example; its state space consists of every vector in $\mathbb{R}^{4}$ with a unit norm, forming the 3-sphere $\mathcal{S}^{3}$. Hence, in such scenarios, by incorporating an operation that enforces these constraints, such as normalization for $\mathcal{S}^{3}$, into the structure of the DNN representing $f^{\mathcal{T}}_{\theta}$, we ensure that transitions remain within the manifold. Moreover, in this way, the distortions that this operation introduces into the dynamical system’s output are factored into the DNN’s optimization process, ensuring they are accounted for.
Alternatively, when we have access to the Riemannian metric of a state space manifold, it is possible to use the exponential and logarithmic maps to do Euclidean calculus in the tangent bundle of the manifold, and map the solution back on the manifold (the reader is referred to [47] for more details).
@@ 5 Experiments
We validate our method with three datasets, each allowing us to study different aspects of it. For evaluation purposes, we use these datasets under the assumption of perfect tracking of the desired state derivatives, $\dot{x}_{t}^{d}$, provided by $f^{\mathcal{T}}_{\theta}$, without any involvement of robots in this process. Subsequently, we test our method in two real-world settings using two different robots. Details on the DNN’s hyperparameters optimization process are presented in Appendix C.
@@ 5.1 Euclidean Datasets
Firstly, we evaluate our method using datasets of Euclidean motions. This enables us to study how different variations of PUMA perform for Euclidean motions and compare the performance of PUMA against state-of-the-art methods for stable Euclidean motion generation.
@@ 5.1.1 LASA
Fig. 7a showcases the accuracy of four variations of PUMA. Here, we compare how different distance metrics in $\ell_{\text{stable}}$ perform for motions in Euclidean spaces, specifically the Euclidean distance and the great-circle (spherical) distance. Furthermore, we also examine the influence of $\ell_{\partial}$ on the accuracy of the learned motions. We use Behavioral Cloning (BC) without a stability loss as a performance upper bound for comparison. Interestingly, each variation of PUMA achieves a similar performance; however, PUMA with a spherical metric shows slightly better performance than PUMA with an Euclidean metric. This suggests that the DNN has no difficulties in mapping the Euclidean space $\mathcal{T}$ into a spherical space $\mathcal{L}$. Lastly, we can also observe that the use of $\ell_{\partial}$ does not harm the accuracy performance of PUMA.
@@
The stability of the motions learned by PUMA hinges on the successful minimization of (8), which we need to empirically test after the learning process concludes. To do this, we integrate the dynamical system over $L$ time steps, starting from $P$ initial states, and observe whether the system converges to the goal. The larger the $P$, the more accurate our results. If $L$ is sufficiently large, the system should reach the goal after $L$ steps. By measuring the distance between the last state visited and the goal, and confirming that it falls below a pre-set threshold $\epsilon$, we can evaluate if a trajectory is successful (i.e., it converges to the goal).
Table 1 provides the stability results of BC and different PUMA variations. The data shows that every variation of PUMA successfully enforces stability across the dataset, generating no unsuccessful trajectories. When compared to BC, which has a 36.4653% rate of unsuccessful trajectories, the benefit of the proposed loss in enforcing stability becomes clear.
@@
Figure 8: Positive invariance evaluation. In figures (a)-(c) white curves represent demonstrations. Red curves represent learned motions when starting from the same initial conditions as the demonstrations. The arrows indicate the vector field of the learned dynamical system.
@@
Fig. 8 demonstrates the effect of the boundary loss $\ell_{\partial}$ in PUMA. Figures 8a and 8b present an example of the qualitative performance of PUMA and CONDOR when no boundary loss is applied. In the top-left region of these images, PUMA exhibits non-smooth trajectories at the boundary, which abruptly change direction due to the applied saturation (see Sec. 4.4). Conversely, CONDOR learns a smoother trajectory in this region. This feature in PUMA disappears when $\ell_{\partial}$ is employed, as depicted in Fig. 8c. Finally, Fig. 8d provides a quantitative evaluation of the boundary loss, confirming our qualitative observations.
In the context of this work, this loss is only relevant for Euclidean state spaces. This is because we do not introduce boundaries in the state space when controlling orientation in $\mathcal{S}^{3}$.
@@ 5.1.2 LAIR
Contrary to the LASA dataset, the LAIR dataset [9] is specifically designed to evaluate second-order motions, i.e., those that map current position and velocity to acceleration. This dataset comprises 10 human handwriting motions, with the state being 4-dimensional, encompassing a 2-dimensional position and velocity. The dataset’s shapes contain multiple position intersections, intentionally designed to necessitate the use of at least second-order systems for their modeling.
@@
The state-of-the-art methods outlined in Sec. 5.1.1, excluding CONDOR, do not address the problem of learning stable second-order systems. This can be attributed to the increased difficulty inherent in learning such systems compared to first-order systems. As a result, we compare the performance of PUMA with that of CONDOR. Fig. 9a illustrates the comparative accuracy of both methods, with PUMA surpassing CONDOR on all metrics. PUMA’s greater learning flexibility allows it to converge to solutions beyond CONDOR’s capabilities. Moreover, this enhanced flexibility, as depicted in Fig. 9b, endows PUMA with a more stable learning process. This is crucial, as motions with unsuccessful trajectories must be discarded. Therefore, if the learned dynamical system is evaluated multiple times throughout a learning process, more learning stability yields a larger set of motions without unsuccessful trajectories, thus offering more alternatives to select from.
@@ 5.2 Non-Euclidean dataset: LASA $\mathcal{S}^{2}$
The LASA $\mathcal{S}^{2}$ dataset [11] comprises 24 motions, each with 3 demonstrations, selected from the LASA dataset. Unlike the LASA dataset, the LASA $\mathcal{S}^{2}$ dataset represents these motions in spherical geometry, as indicated by its name, in $\mathcal{S}^{2}$. The sphere, where the motions evolve, is structured similarly to unit quaternions, though with one less dimension. Essentially, we have a 3-dimensional Euclidean space where vectors are constrained to have unit norm. As a result, the state space is a 2-dimensional spherical manifold embedded in this 3-dimensional space. This setup allows us to examine the performance of PUMA in a manifold with similar attributes to that of unit quaternions. However, the simpler visualization of this setup facilitates a more intuitive analysis of PUMA’s performance.

### v3

@@ 4.2.3 Loss Function
Crucially, the surrogate stability conditions from Theorem 4 can be enforced in a DNN by minimizing the following expression, $\forall y_{0}\in\mathcal{L}$ and $\forall t\in\mathbb{R}_{\geq 0}$,
@@
where $m\in\mathbb{R}_{>0}$. This function resembles the form of the triplet loss introduced in Sec. 3.4. However, in our context, $y_{\mathrm{g}}$ serves as the anchor sample, $y_{t+\Delta t}$ as the positive sample, and $y_{t}$ as the negative sample. In the following discussion, we explore how this expression induces the fulfillment of the surrogate stability conditions within a DNN.
@@
Figure 5: Left: Illustration of the effect of optimizing $\ell_{\text{stable}}$. The red arrow depicts $y_{t}$ and $y_{t+\Delta t}$ being modified to fulfill (17). Right: Example of trajectories generated with $f^{\mathcal{T}}_{\theta}$ and $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$ post-training. Each $y_{t+\Delta t}$ is closer to $y_{\mathrm{g}}$ than its predecessor $y_{t}$.
For any $y_{0}$ where $x_{0}\neq x_{\mathrm{g}}$, minimizing (16) enforces transitions to progressively approach $y_{\mathrm{g}}$, as its minimization implies
$d(y_{\mathrm{g}},y_{t+\Delta t})+m\leq d(y_{\mathrm{g}},y_{t}).$ (17)
Hence, $d(y_{\mathrm{g}},y_{t+\Delta t})<d(y_{\mathrm{g}},y_{t})$, i.e., the second surrogate condition (see Fig. 5).
@@
In the case where $y_{0}=y_{\mathrm{g}}=\psi_{\theta}(x_{\mathrm{g}})$, that is, when the system is initialized at the equilibrium, enforcing a value of $y_{t+\Delta t}$ different from $y_{\mathrm{g}}$ would only increase the function in (16). Consequently, for $y_{0}=y_{\mathrm{g}}=\psi_{\theta}(x_{\mathrm{g}})$, (16) achieves its minimum when $y_{t+\Delta t}=y_{\mathrm{g}}=y_{0}$, equal to $m$. Therefore, this equation also enforces the first surrogate condition.
@@
In this equation, $\mathcal{B}^{s}$ denotes a batch of initial latent states $y_{0}$. These states are derived by mapping initial states $x_{0}$ (sampled randomly from $\mathcal{T}$) via the function $\psi_{\theta}$. Meanwhile, $\mathcal{H}^{s}$ represents a set of time instants $t$ at which the loss is minimized.
@@ 4.3 On the Stability Loss Metric
To elaborate, recall that the learned dynamical system can be expressed as $\dot{x}_{t}=\phi_{\theta}(y_{t})$. It then follows that the output of our model is entirely determined by the latent state $y_{t}$. This implies that if two different states $x^{\text{a}}_{t}$ and $x^{\text{b}}_{t}$ map to the same latent state, that is, $y_{t}=\psi_{\theta}(x^{\text{a}}_{t})=\psi_{\theta}(x^{\text{b}}_{t})$, the time derivative computed by the learned dynamical system for both states will be identical. Such behavior would manifest in certain states if $\psi_{\theta}$ were a non-bijective [^4] More rigorously, the property being described is that of non-injective functions. However, for practical purposes, we can define the codomain of $\psi_{\theta}$ to be equal to its image, which is the property that an injective function must have to be bijective. Hence, in this specific context, we use these terms interchangeably. function. It would, therefore, be potentially harmful for the learning process to enforce $\psi_{\theta}$ to be non-bijective, as this would constrain the family of solutions to which the system can converge, hindering the DNN’s optimization process. Ideally, we would like $\psi_{\theta}$ to have the capacity to converge to a bijective function if required, where each state $x_{t}$ maps to a unique latent state $y_{t}$. Consequently, for any state $x_{t}$, it would be possible to compute a value of $\dot{x}_{t}$ that is independent of those calculated for any other state. Fig. 6 presents an example of bijective and non-bijective mappings between dynamical systems where the surrogate stability conditions are enforced.
@@ 4.3.1 Homeomorphisms and State Space Geometry
In our context, this implies that the distance function employed in the stability loss (18) must induce a topology in $\mathcal{L}$ that is equivalent to that of $\mathcal{T}$. For example, if orientations are described using unit quaternions, the topology of $\mathcal{T}$ would be spherical. Then, the stability loss metric should generate a topology that is homeomorphic to the sphere. Otherwise, it would be infeasible for the DNN to establish a homeomorphism between $\mathcal{T}$ and $\mathcal{L}$.
In this work, Sec. 5, we use unit quaternions to represent orientations in our robot experiments. For a detailed discussion on metrics and pose control in this context, the reader is referred to Appendix B.
@@ 4.4 Boundary Conditions
Lastly, it is crucial to ensure that a dynamical system evolving in $\mathcal{T}$ always remains within this manifold. Two scenarios are relevant to this work: 1) ensuring $\mathcal{T}$ is positively invariant with respect to $f^{\mathcal{T}}_{\theta}$; and 2) considering the state’s geometry when computing the evolution of the dynamical system.
@@ 4.4.1 Positively Invariant Sets
To address this, we design the dynamical system so that it cannot leave $\mathcal{T}$ by construction. This can be achieved by projecting any transitions that would leave $\mathcal{T}$ back onto its boundary. For example, in Euclidean state spaces, $\mathcal{T}$ can be represented as a hypercube; therefore, in this case, this projection is achieved by saturating/clipping the points that leave $\mathcal{T}$. Furthermore, we found that introducing an additional loss, denoted as $\ell_{\partial}$, can be beneficial in enforcing this condition through the optimization process of the DNN. As proposed in [6], this can be accomplished by using the scalar product between the dynamical system’s velocity $v(x_{t})$ (which equates to $f^{\mathcal{T}}_{\theta}$ for first-order systems) and the outward-pointing normal vector $n(x_{t})$ at states within the boundary of $\mathcal{T}$. This product should be ensured to be equal to or less than zero. Then, to achieve this for DNNs, we introduce the following loss function
@@
4.4.2 Evolving in Non-Euclidean State Spaces
@@
Non-Euclidean state spaces are commonly defined within a higher-dimensional Euclidean space and can sometimes be constructed by incorporating constraints into this space. Consider the unit quaternion as an example: its state space consists of every vector in $\mathbb{R}^{4}$ with a unit norm, forming the 3-sphere $\mathcal{S}^{3}\subseteq\mathbb{R}^{4}$. Therefore, in such scenarios, by integrating an operation that enforces these constraints, such as normalization for $\mathcal{S}^{3}$, into the structure of the DNN representing $f^{\mathcal{T}}_{\theta}$, we ensure that transitions remain within the manifold. Moreover, in this way, the distortions that this operation introduces into the dynamical system’s output are factored into the DNN’s optimization process, ensuring they are accounted for.
Alternatively, when we have access to the Riemannian metric of a state space manifold, it is possible to use the exponential and logarithmic maps to do Euclidean calculus in the tangent bundle of the manifold, and then map the solution back on the manifold (the reader is referred to [49] for more details).
@@ 5 Experiments
We validate our method with three datasets, each allowing us to study different aspects of it. For evaluation purposes, we use these datasets under the assumption of perfect tracking of the desired state derivatives, $\dot{x}_{t}^{d}$, provided by $f^{\mathcal{T}}_{\theta}$, without any involvement of robots in this process. Subsequently, we test our method in two real-world settings using two different robots. We have made our code implementation of PUMA publicly available at: https://github.com/rperezdattari/Deep-Metric-IL-for-Stable-Motion-Primitives. Details on the DNN’s hyperparameters optimization process are presented in Appendix C.
@@ 5.1 Euclidean Datasets
First, we evaluate our method using datasets of Euclidean motions. This approach allows us to examine the performance of different variations of PUMA for Euclidean dynamical systems and compare their effectiveness against state-of-the-art methods.
@@ 5.1.1 LASA
Fig. 7a showcases the accuracy of four variations of PUMA. Here, we compare the performance of different distance metrics in $\ell_{\text{stable}}$ for motions in Euclidean spaces, focusing on the Euclidean distance and the great-circle (spherical) distance. Furthermore, we also examine the influence of $\ell_{\partial}$ on the accuracy of the learned motions. We use Behavioral Cloning (BC) without a stability loss as an upper performance bound for comparison. Interestingly, each variation of PUMA achieves a similar performance; however, PUMA with a spherical metric shows slightly better performance than PUMA with an Euclidean metric. This suggests that the DNN has no difficulties in mapping the Euclidean space $\mathcal{T}$ into a spherical space $\mathcal{L}$. Lastly, we can also observe that the use of $\ell_{\partial}$ does not harm the accuracy performance of PUMA.
@@
The stability of the motions learned by PUMA hinges on the successful minimization of (8), which we need to test after the learning process concludes empirically. To do this, we integrate the dynamical system over $L$ time steps, starting from $P$ initial states, and observe whether the system converges to the goal. The larger the $P$, the more accurate our results. If $L$ is sufficiently large, the system should reach the goal after $L$ steps. By measuring the distance between the last state visited and the goal, and confirming that it falls below a pre-set threshold $\epsilon$, we can evaluate if a trajectory is successful (i.e., it converges to the goal).
Table 1 provides the stability results of BC and different PUMA variations. The data shows that every variation of PUMA successfully enforces stability across the dataset, generating no unsuccessful trajectories. Compared to BC, which has a 36.4653% rate of unsuccessful trajectories, the benefit of the proposed loss in enforcing stability becomes clear.
@@
Figure 8: Positive invariance evaluation. In figures (a)-(c), white curves represent demonstrations. Red curves represent learned motions when starting from the same initial conditions as the demonstrations. The arrows indicate the vector field of the learned dynamical system. Figure (d) provides a positive invariance comparison between CONDOR and the variations of PUMA.
@@
Fig. 8 demonstrates the effect of the boundary loss $\ell_{\partial}$ in PUMA. Figures 8a and 8b present an example of the qualitative performance of PUMA and CONDOR when no boundary loss is applied. In the top-left region of these images, PUMA exhibits non-smooth trajectories at the boundary, which abruptly change direction due to the applied saturation (see Sec. 4.4). Conversely, CONDOR learns a smoother trajectory in this region. This feature in PUMA vanishes when $\ell_{\partial}$ is applied, as depicted in Fig. 8c. Finally, Fig. 8d provides a quantitative evaluation of the boundary loss, confirming our qualitative observations.
In this work, this loss is only relevant for Euclidean state spaces because we do not introduce boundaries in the state space when controlling orientation in $\mathcal{S}^{3}$.
@@ 5.1.2 LAIR
Contrary to the LASA dataset, the LAIR dataset [9] is specifically designed to evaluate second-order motions, i.e., those that map current position and velocity to acceleration. This dataset comprises 10 human handwriting motions, with the state being 4-dimensional, encompassing a 2-dimensional position and velocity. The dataset’s shapes contain multiple position intersections, intentionally designed to necessitate the use of at least second-order systems for their successful modeling.
@@
The state-of-the-art methods outlined in Sec. 5.1.1, excluding CONDOR, do not address the challenge of learning stable second-order systems. This can be attributed to the greater difficulty inherent in learning such systems compared to first-order systems. As a result, we compare the performance of PUMA with that of CONDOR. Fig. 9a illustrates the comparative accuracy of both methods, with PUMA surpassing CONDOR on all metrics. PUMA’s greater learning flexibility allows it to converge to solutions beyond CONDOR’s capabilities. Moreover, this enhanced flexibility, as depicted in Fig. 9b, endows PUMA with a more stable learning process. In practice, this makes a significant difference, as motions with unsuccessful trajectories must be discarded when considering the system’s stability. Therefore, greater learning stability results in a larger set of motions without unsuccessful trajectories, providing more alternatives of successfully learned systems to choose from.
@@ 5.2 Non-Euclidean dataset: LASA $\mathcal{S}^{2}$
The LASA $\mathcal{S}^{2}$ dataset [11] comprises 24 motions, each with 3 demonstrations. Unlike the LASA dataset, the LASA $\mathcal{S}^{2}$ dataset represents these motions in spherical geometry, as indicated by its name, in $\mathcal{S}^{2}$. The sphere, where the motions evolve, is structured similarly to unit quaternions, though with one less dimension. Essentially, we have a 3-dimensional Euclidean space where vectors are constrained to have a unit norm. As a result, the state space is a 2-dimensional spherical manifold embedded in this 3-dimensional space. This setup allows us to examine the performance of PUMA in a manifold with similar attributes to those of unit quaternions. However, the more straightforward visualization of this setup facilitates a more intuitive analysis of PUMA’s performance.